OpenStudy (anonymous):

Suppose f(x)= ln3x and f^-1X denotes the inverse of f. Then the integral of f^-1x dx =

OpenStudy (anonymous):

my work so far: x= ln (3y) e^x= 3y y = E^x /3 (that would be e to the x, all divided by 3) y= 1/3 e^x this is where i get confused. i know the antiderivative of e^x is e^x, but i dont know how it works with a constant in front. my textbook isnt helping at all!